A spherical balloon is deflated so that its volume is decreasing at a rate of 5 ftଷ/min. How
fast is the diameter of the balloon decreasing when the radius is 3 ft?
1
Expert's answer
2020-06-09T13:17:57-0400

As per the given question,

Rate of the change in the volume of the balloons"(\frac{dV}{dt})=5 ft^3/min"

Final diameter of the volume "= 3ft"

We know that volume of the sphere "(V)=\frac{4 \pi r^3}{3}"


"\Rightarrow \frac{dV}{dt}=4\pi3r^2 \frac{dr}{3dt}=4\pi r^2\frac{dr}{dt}"


"\Rightarrow 5ft^3/min =4\pi \times 3^2 \frac{dr}{dt}"

"\Rightarrow \frac{dr}{dt}=\frac{5}{4 \pi \times 9}=\frac{5}{36 \pi}ft/sec"